Total Materials Variance

Young Inc. produces plastic bottles. Production of 16-ounce bottles has a standard unit quantity of 0.45 ounce of plastic per bottle. During the month of June, 240,000 bottles were produced using 110,000 ounces of plastic. The actual cost of plastic was $0.042 per ounce, and the standard price was $0.045 per ounce. There is no beginning or ending inventories of plastic.

Required:

Calculate the total variance for plastic for the month of June. Enter the amount as a positive number and select Favorable or Unfavorable.
$ ———— Favorable
Answer #1

Materials price variance = 110000*(0.045-0.042) = 330 Favourable

Materials usage variance =((240000*0.45)-110000)*0.045 = -90 Unfavorable

Total variance =330-90 =240 Favorable